# Transformer Block Coupling and its Correlation with Generalization in LLMs

### v2

@@ 5.3 Emergence of Coupling with Training
Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure LABEL:fig:paris_all_8to16 ) and across tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout). Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 12B, 6.9B, and 1B (Biderman et al., 2023) (Figure 2), and observe that coupling is generally low at initialization and increases significantly during early training steps.
@@ 5.5 Regularity in Hidden Trajectories
We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs. The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure 13). Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training). Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
Figure 13: Linearity of Trajectories. The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 5.5), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. Models are sorted by number of parameters.
5.6 Layer-Wise Exponential Growth
Across all LLMs considered, the hidden trajectories exhibit exponential growth that emerges with training (Figure 14). The majority of the models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. The expodistance measurements show little dependence on model parameters, and take on consistently low values across all featured LLMs (Figure 14). In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 22). Under certain assumptions, exponential spacing is justified in Section 6.2.
Figure 14: Coefficient of variation of layer-wise exponent. The figure presents the coefficient of variation of the layer-wise exponent (Section 3.3) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for trained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale.
@@ 6 Discussion
6.1 Correlation with Generalization
As discussed in Section 5.4, stronger coupling across Jacobian singular vectors is related to improved generalization (Figure 1). This relationship may assist in evaluating the training of transformers. In particular, this finding suggests that the development of training methods to amplify coupling across transformer blocks may lead to favourable model performance.
6.2 Emergence of Exponential Spacing with Coupling
The emergence of exponential spacing in many LLMs can be viewed as a result of coupling under some assumptions on the homogeneity of the singular values across depth. Strong coupling suggests that the linearizations of hidden trajectories satisfy a linear difference equation (see Appendix A.4), given by
$x^{l+1}-x^{l}=Ax^{l}$
Considering the singular value decomposition of $A$ in Equation 18 as $A=U\Sigma V^{T}$, and assuming that there is coupling ($U=V$), the difference equation can be diagonalized in the same basis across depth. Further assuming that the top singular values of the Jacobian blocks display low variation across depth, (Figure LABEL:fig:svalues_trained), we expect
$||x^{l}||\approx e^{\sigma l}\,||x^{0}||,$ (16)
where $\sigma$ is the top singular value. Since we assume $\sigma$ is independent of depth, the above norm is expected to be exponential in $l$, and offers an explanation to the measurements in Figure 14.
@@ 7 Related Work
Residual Networks. ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021). The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment. Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023). We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023). Our measurements quantify the emergence with training of linearity and spacing, including at different depths, which was previously observed mostly qualitatively. Further comparison is included in Appendix A.3.
@@
Neural ODE. Introduced by (Chen et al., 2018), Neural ODE view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE. (Zhong et al., 2022; Li et al., 2021) studied neural ODE specific perspective to transformers and other LLM architectures. In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
Information Processing in LLMs. Previously, Katz & Belinkov (2023); Bietti et al. (2023) have described memory and semantic flow in intermediate states of transformers while van Aken et al. (2019) have examined BERT hidden trajectories. Investigation of LLMs for reasoning tasks and interpretability remains an important research focus (Huang & Chang, 2023).
@@
In-context learning. Previous works have observed that LLMs are able to perform in context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024): the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context. Their results consider a single layer of trained self-attention as implementing a single step of a gradient-descent-like learning rules across transformers depth, minimizing the MSE of a linear model
$x_{\text{final}}=\min_{x}||Ax-b||^{2}$
A gradient-descent learning rule would cause depth-wise updates to representations of the form
$x_{t+1}=x_{t}+\epsilon A^{T}(Ax_{t}-b)=(I+\epsilon A^{T}A)x_{t}-{\epsilon}A^{T}b$
We should expect similarity in the matrices $I+{\epsilon}A^{T}A$ across depth, which is suggested by coupling.
Hidden Representation Dynamics. Previous studies (Gai & Zhang, 2021; Haber & Ruthotto, 2017; Ee, 2017), have focused on interpreting deep neural networks from a dynamical systems perspective. Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space (Gai & Zhang, 2021). Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts (Nam et al., 2023). For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
@@
Structure in Hidden Representations. Emergence of regularity in last-layer representations during training was noticed in Neural Collapse (Papyan et al., 2020). Subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023). Within LLMs, recent works have identified uniform structure in token means (Wu & Papyan, 2024), and low dimensional structure across hidden trajectories (Sarfati et al., 2024). Our work studies the representations of all layers in LLMs in a local sense through the interactions of tokens through their Jacobians.
@@ 8 Limitations
A limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them. Several of these models are trained independently of each other, without carefully controlling for different experimental parameters. Consequently, direct comparisons between models become challenging, and so determining the cause of varying levels of regularity between LLMs becomes difficult. Nonetheless, this makes the consistent emergence of these properties interesting for further study.
@@ 9 Conclusion
Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions. Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations. We additionally observe and quantify the emergence of linearity and exponential spacing with training in the embedding trajectories of LLMs. Notably, we establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization. These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
@@ A.3 Comparison to ResNets
Linearity and Equidistance. Linearity in hidden trajectories, as observed in ResNets Li & Papyan (2023); Gai & Zhang (2021), also emerges with training in LLMs. The mean LSS value among the evaluated LLMs is $4.24$ (Figure 13), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension). In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories. In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 14). We quantify this spacing through a low coefficient of variation, displaying the presence of exponential growth in token trajectories. In both classifier ResNets and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
@@
A.4 Dynamical Motivation
The equality of top left and right singular vectors suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
$x^{\prime}=Ax$ (17)
Its solution is given by
$x(t)=\exp(tA)x(0)=\lim_{L\to\infty}\left(I+\frac{At}{L}\right)^{L}x(0)$ (18)
where the product is repeated $L$ times. Expanding the brackets shows that $x(t)$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. (2016) which views ResNets as
$x(t)=\left(I+\frac{tA_{1}}{L}\right)\left(I+\frac{tA_{2}}{L}\right)\dots\left(I+\frac{tA_{L}}{L}\right)x(0)$ (19)
However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{i}$ matrices. The coupling phenomenon suggests ResNets as implementing the simpler system
$x(t)=\left(I+\frac{tA}{L}\right)\left(I+\frac{tA}{L}\right)\dots\left(I+\frac{tA}{L}\right)x(0),$ (20)
where all the $A_{i}$ matrices are aligned. One benefit of this interpretation is that, as $L\to\infty$, we can write $x(t)$ in a simple closed form, namely $x(t)=\exp(tA)x(0)$.
Borrowing from the dynamical perspective, we quantify the similarity of hidden trajectories to the evolution of a linear ODE, in order to detect the emergence of a simple linearization to representations.
A.5 Experimental and Implementation Details

### v3

@@ 5.3 Emergence of Coupling with Training
Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure 3 ) and across tokens (Figures 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout). Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 6.9B and 12B (Biderman et al., 2023) (Figure 1b), and observe that coupling is generally low at initialization and increases persistently throughout training. Moreover, there is a clear sense of locality in the strength of coupling which is visually displayed in Figure (1c-e)
The gradual growth of coupling observed in Figure (1b) parallels the logarithmic increase in accuracy during training for Pythia 12B and Pythia 6.9B (Biderman et al., 2023). This highlights the relationship between coupling and performance, since both properties emerge at similar training iterations and rates.
@@ 5.4 Correlation with Generalization
The results for ViTs demonstrate that stochastic depth encourages coupling during training (Figure LABEL:figure:coupling_vitb) and that coupling correlates with accuracy when fixing SD rate (Figure LABEL:figure:coupling_vita). This finding suggests that coupling may provide new insight into stochastic depth’s underlying mechanism, and that developing training methods to amplify coupling across transformer blocks could provide additional regularization and improve performance.
We hypothesize that simple trajectories may lead to better generalization. This agrees with many generalization bounds in machine learning (Arora et al., 2018), which suggest that models with lower complexity tend towards better generalization. Additionally, prior works (Novak et al., 2018) demonstrate that the Frobenius norm of input-output Jacobians is related to generalization, providing evidence that coupling — a structural property derived from Jacobians — may also correlate with generalization.
@@ 5.5 Regularity in Hidden Trajectories
We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs. The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure LABEL:figs:lss-expo). Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training). Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, MPT 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
Among the LLMs considered, the vast majority of hidden trajectories exhibit exponential growth that emerges with training (Figure LABEL:fig:expo-all b). Many models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 23). Under certain assumptions, exponential spacing is motivated in Section 6.1.
@@ 6 Discussion
6.1 Emergence of Regularity with Coupling
Under certain assumptions, the emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling. Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$
We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. The representations follow the linearized equation:
$x^{l+1}=x^{l}+J_{l}x^{l}=(I+J_{l})x^{l}.$
Expanding across layers, the entire system can be approximated by the product
$x^{L}=(I+J_{L})(I+J_{L-1})\cdots(I+J_{1})x^{0}.$
Assuming that $J_{l}\approx USU^{T}$, this equation predicts that the norm of $x_{l}$ would exhibit exponential growth layer by layer. Expanding $U$ and $S$,
$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}.$
where $u_{j}$ and $s_{j}$ represent the eigenvectors and eigenvalues of the Jacobian, respectively. In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J$. However, in our experiments, we observe a notable tendency towards linearity, suggesting that the representations align progressively during training with the eigenvectors of the coupled Jacobians.
6.2 Significance of Coupling
The coupling phenomenon provides insight into the internal operations of transformer. We hypothesize that during training, the LLM learns to represent embeddings in specific low-dimensional subspaces (Eldar & Mishali, 2009). Given an input, the first layer converts the input into embeddings within one of these learned subspaces. Each subsequent transformer layer modifies these embeddings, potentially moving them to different subspaces. Strong coupling between consecutive layers suggests that the LLM tends towards representations in the same or similar subspaces across many layers. Weak coupling suggests that the subspaces may change between layers, though usually gradually, and that adjacent layers still operate in relatively similar subspaces. Previous works have shown (Lad et al., 2024; Gromov et al., 2024) that the early and late layers of language models behave differently, which may be understood through coupling; tokens remain in similar spans, then transition to a different subspace, continuing within a new span that is consistent in the remainder of the transformer.
The emergence of coupling with training steps (Figure 1) may provide insight into the dynamics. Under full coupling and a difference equation approximation, the representations evolve as
$x^{l}=\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{l}u_{j}^{\top}x^{0}$
where $u_{i}$ and $\lambda_{i}$ denote the eigenvectors and eigenvalues of the Jacobian, respectively. In this case, the gradients of the loss $L$ with respect to prediction $y$, $x^{L}$ are represented by
$\frac{\partial\text{L}}{\partial x^{0}}(x^{L},y)\approx\sum_{j=1}^{d_{\text{model}}}u_{j}(1+s_{j})^{L}u_{j}^{\top}(y-x^{L}),$
Due to the coupling, the dynamics exhibit either exponential growth or decay in different subspaces, depending on the sign of $s_{j}$, which is known to cause challenges for optimization as in past works on dynamical isometry (Pennington et al., 2017). We infer that increasing coupling during training makes optimization progressively more difficult. Conversely, as training progresses, it becomes harder to achieve stronger coupling, and is consistent with the logarithmic trend in Figure 1.
@@ 7 Related Work
Residual Networks. ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021). The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment. Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023). We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023). Further comparison is included in Appendix A.3.
@@
Neural Ordinary Differential Equations. Neural ODEs (Chen et al., 2018) view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE, with some extensions to transformers (Zhong et al., 2022; Li et al., 2021) The emergence of coupling in transformers suggests that a discretization of a simple itertaive process emerges in LLMs.
@@
In-context learning. LLMs can perform tasks through examples provided in a single prompt, demonstrating in-context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024). Studies suggest trained self-attention layers implement gradient-descent-like updates across depth to minimize the MSE of a linear model:
$x_{\text{final}}=\min_{x}||Ax-b||^{2}.$
These updates take the form:
$x_{t+1}=(I+\epsilon A^{T}A)x_{t}-\epsilon A^{T}b.$
Coupling across depth suggests similarity in the matrices $I+\epsilon A^{T}A$.
Hidden Representation Dynamics. Prior research interprets deep neural networks through dynamical systems, revealing that training trajectories align with geodesic curves (Gai & Zhang, 2021) and partition activation space into basins of attraction (Nam et al., 2023). For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
@@
Structure in Hidden Representations. Neural Collapse (Papyan et al., 2020) highlights emergent regularity in last-layer representations, with subsequent studies exploring hidden-layer structures and their theoretical underpinnings (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023). In LLMs, recent works identify uniform token structures (Wu & Papyan, 2024) and low-dimensional hidden trajectories (Sarfati et al., 2024). Our work examines local token interactions through Jacobian dynamics across all LLM layers.
@@ 8 Limitations
Our analysis is limited to pretrained LLMs and their fine-tuned variants due to the high computational cost of training large models. Variations in experimental setups across independently trained models hinder direct comparisons, making it challenging to pinpoint causes of differing regularity. However, the consistent emergence of these properties warrants further study.
@@ 9 Conclusion
Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions. Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations. We establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization. These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
@@ A.3 Comparison to ResNets
Linearity and Equidistance. Linearity in hidden trajectories, as observed in ResNets Li & Papyan (2023); Gai & Zhang (2021), also emerges with training in LLMs. The mean LSS value among the evaluated LLMs is $4.24$ (Figure LABEL:fig:lss-all), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension). In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories. In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure LABEL:fig:expo-all). We quantify this spacing through a low coefficient of variation, displaying the presence of exponential growth in token trajectories. In both classifier ResNets and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
@@
A.4 Significance of Coupling
The transformer block coupling phenomenon offers insight into several prominent practices in LLM research, as summarized in Table 2.
Table 2: Significance of the Coupling Phenomenon. A table which highlights the implications of transformer block coupling to a variety of effors in machine learning research.
Current Research Practice Key Idea Our Contribution
Compressing models by merging blocks (Fu et al., 2022; Kim et al., 2024) Combine adjacent transformer blocks to reduce model size without significant performance loss. Demonstrates that merging is effective because blocks become strongly coupled during training.
Compressing models by pruning blocks (Elkerdawy et al., 2020; Kim et al., 2024; Dror et al., 2021; Fang et al., 2023) Remove certain transformer blocks while preserving functionality. Explains that pruning works because the coupling ensures redundancy across blocks.
Compressing models by projecting weight matrices (Ashkboos et al., 2024) Reduce dimensionality by projecting weights into smaller subspaces. Shows that coupling induces a low-dimensional subspace in which blocks’ weights are aligned.
Studying the effect of transformer block permutations (Hu et al., 2021; Mahabadi et al., 2021; van der Ouderaa et al., 2024; Li et al., 2016) Investigate whether permuting the order of blocks affects model performance. Explains why permutations have minimal impact: strong coupling creates structural robustness.
Early exiting in LLMs (Scardapane et al., 2020; Jazbec et al., 2024) Allow models to exit computation early based on task confidence. Reveals that early exiting works because representations progress linearly along a shared trajectory due to coupling.
A.5 Dynamical Motivation
The equality of top left and right singular vectors suggests that the linearizations form a simple linearized system that acts on representations. Consider a difference equation
$x^{l+1}-x^{l}=A_{l}\,x^{l}$ (19)
Its solution at the final $L$ is given by
$x^{L}=\prod_{l=1}^{L}(I+A_{l})\,x^{0}$ (20)
Expanding the brackets shows that $x^{l}$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. (2016) which views ResNets as
$x^{l}=\left(I+A_{l-1}\right)\left(I+A_{l-2}\right)\dots\left(I+A_{1}\right)x^{0}$ (21)
However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{l}$ matrices. The coupling phenomenon suggests the model as implementing the simpler system
$x^{l}=\left(I+A\right)^{l}x^{0}$ (22)
where all the $A$ matrices are aligned. One benefit of this interpretation is that, we can write $x^{l}$ in a simple closed form, as above. We quantify the similarity of hidden trajectories to the evolution of the above difference equation, in order to detect the emergence of a simple linearization to representations. The emergence of increased linearity and exponential spacing in many LLMs can be analyzed as a result of coupling under some conditions on the spectral decomposition of the Jacobians. Considering input embeddings $x_{1}^{0},\ldots,x_{n}^{0}$ and the linearization of the last token embedding $x_{n}^{l}$ given by $J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})$:
$x_{n}^{l}=(I+J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0}))x_{n}^{l-1}$ (23)
We simplify notation and write $x_{n}^{l}=x^{l},\,J_{n,n}^{l}(x_{1}^{0},\ldots,x_{n}^{0})=J^{l}$. Under the assumption of spectral coupling, $J^{l}=U_{l}S_{l}V_{l}^{T}\approx US_{l}U^{T}$, and the linearized effect of the last token is
$x^{L}=\prod_{l=1}^{L}(I+U_{l}S_{l}V_{l}^{T})x^{0}\approx U\left(\prod_{l=1}^{L}(I+S_{l})\right)U^{T}x^{0}$ (24)
Suppose that $x^{0}=u_{k}$ is the $k$-th left singular vector of $J^{l}$. It follows that $x^{L}=\prod_{l=1}^{L}(1+s_{k}^{l})x^{0}$, where $s_{k}^{l}$ denotes the $k$-th singular value at layer $l$. The exponential spacing measurement is motivated by the consistent choice $s_{k}=s_{k}^{1}=s_{k}^{2}=\cdots=s_{k}^{L}$. Explicitly, $x^{L}=(1+s_{k})^{L}x^{0}$, and by Equation 17, for each $l$
$\alpha_{k}^{l}=\ln\left(\frac{(1+s_{k})^{l}||u_{0}||}{(1+s_{k})^{l-1}||u_{0}||}\right)=\ln(1+s_{k})\implies\text{ED}=0$
that is, the coefficient of variation 0 across $l$. In addition, if $x^{l}=(1+s_{k})x^{l-1}$, it is clear that the trajectory would form a perfect line, yielding $LSS=1$ by the discussion in Section 3.2. In general, trajectories are not expected to be perfectly linear unless $x^{0}$ aligns with an eigenvector of $J^{l}$.
A.6 LLM Evaluation and Implementation Details